You own a bond that is callable in 3.5 years. The bond has a 7% coupon, paying interest semiannually, with a par value of $1000. the yield-to-call is 5.6%. what is the call premium if the bond currently sells for $ 1109.88?

A 60

B 70

C 80

D 90

E 100
Answer #1
Price of the bond is the sum of PVs of the:
*Call value and
*Semi annual interest
PV of semi annual interest = 35*(1.028^7-1)/(0.028*1.028^7) = $         219.71
PV of call value = 1109.88-219.71 = $         890.17
Therefore call value = 890.17*1.028^7 = $     1,080.00
Call premium = 1080-1000 = $           80.00
Answer: Option [C] $80